Break-even analysis for a service company

Sprint Corporation (S) is one of the largest digital wireless service providers in the United States. In a recent year, it had approximately 60 million direct subscribers (accounts) that generated revenue of $33,347 million. Costs and expenses for the year were as follows (in millions):

Cost of revenue $14,958
Selling, general, and administrative expenses 7,994
Depreciation and amortization 8,150

Assume that 30% of the cost of revenue and 70% of the selling, general, and administrative expenses are fixed to the number of direct subscribers (accounts). In part (a) and (b), round all interim calculations and final answers to one decimal place.

a. What is Sprint’s break-even number of accounts, using the data and assumptions given?
million accounts

b. How much revenue per account would be sufficient for Sprint to break even if the number of accounts remained constant?
$ million per account
Answer #1

1. Sprint’s break-even number of accounts, using the data and assumptions given

% of fixed cost Fixed cost
cost of revenue 14958 30% 4487
selling general and administrative 7994 70% 5596
depreciation 8150 100% 8150
Total fixed cost 18233$

2. Revenue per account

=33347/60

556

3.variable cost per unit

% of fixed cost Fixed cost
cost of revenue 14958 70% 10471
selling general and administrative 7994 30% 2398
depreciation 8150 0% 0
Total variable cost 12869$

variable cost per unit = 12869/60account

=214$

4. Break even accounts

=fixed cost / contribution margin per unit

=18233 / (revenue per account-variable cost per account)

=18233/(556-214)

=18233/342

=53.3 accounts

b)Suppose price is X

break even unit s= 60 mil

60 = 18233/(x-214)

60x-12840=18233

x=517.9 $